Parantheses in equations:
I have tried now about 6 times working on these problems and I just can't get it all together. I get one part then mess up another, please HELP.
Problem #1
-(7D - 8) = -5(2B - 3) +1 I need to solve and do the check.
This is what I have done.
-(7D- 8) = -5(2D - 3) +1
-7D + 8 = -10D + 15 + 1
-7D + 8 = -10D + 16
+7D +7D
+8 = -3D + 16
-16 -16
-8 = -3D
D = -8 / -3
CHECK:
-(7D - 8) = -5(2D - 3) + 1
-(7 (-8 / -3) -8) = -5 (2 (-8 / -3) -3) +1
-(56 / -3 -8) = -5 (-16 / -3) -3) +1
Please help

Answer by venugopalramana(3286)

SEE MY COMMENTS BELOW..
GOOD TO SEE YOU TRY SINCERELY.KEEP IT UP! THAT IS THE BEST WAY OF LEARNING.
--------------------------------------------------------------
Parantheses in equations:
I have tried now about 6 times working on these problems and I just can't get it all together. I get one part then mess up another, please HELP.
Problem #1
-(7D - 8) = -5(2B - 3) +1 I need to solve and do the check.
This is what I have done.
-(7D- 8) = -5(2D - 3) +1..........OK
-7D + 8 = -10D + 15 + 1..........VERY GOOD
-7D + 8 = -10D + 16..............VERY GOOD
+7D +7D
+8 = -3D + 16...........NO
TRY TO BRING UNKNOWNS TO LHS...AND KNOWNS TO RHS...
TO DO THIS FOLLOW THE RULE THAT WHEN YOU CHANGE SIDES FROM LEFT TO RIGHT OR OTHERWISE..+ WILL BECOME - AND - WILL BECOME + ON THE OTHER SIDE.
1.WE HAVE TO SEND -10D AN UNKNOWN ON RHS FROM RHS TO LHS. SO IT WILL BECOME +10D ON LHS
2.WE HAVE TO SEND +8 A KNOWN ON LHS FROM LHS TO RHS.SO IT WILL BECOME -8 ON RHS.
SO WE GET
-7D+10D = 16 - 8 =8
3D = 8
D = 8/3
------------------------------------------------
CHECK
-(7D-8) = -5(2B-3)+1
LHS = -[7*8/3 -8]= -[(7*8-8*3)/3]=-(56-24)/3=-32/3
RHS = -5(2*8/3 - 3) + 1 = -5[(2*8 - 3*3)/3]+1 = -5*7/3 +1
= (-35+3)/3=-32/3 = LHS..OK..
-16 -16
-8 = -3D
D = -8 / -3
CHECK:
-(7D - 8) = -5(2D - 3) + 1
-(7 (-8 / -3) -8) = -5 (2 (-8 / -3) -3) +1
-(56 / -3 -8) = -5 (-16 / -3) -3) +1
Please help